# A Mathematical Framework for Misinformation Propagation in Complex Networks: Topology-Dependent Distortion and Control

### v2

@@ I Introduction
Quantification and dynamics of misinformation have been studied across journalism, sociology, neuroscience, collective biological systems, distributed computing, and networked cyber-physical systems, but a unified mathematical understanding remains limited. Prior studies have offered domain-specific or simulation-driven insights, such as qualitative analyses of error propagation based on connectivity measures Krol_Kukla_2012, error accumulation in deep neural architectures Li2017, and preferred interaction topologies in sociotechnical and biological settings klemm2005. Efforts to restrict misinformation on online platforms Ghoshal2025 and to characterize the spread of true versus false information in complex networks NIAN2021 provide valuable perspectives, but do not yield a general analytical framework for quantifying misinformation in distributed systems.
Information flow in any multi-agent system is fundamentally shaped by connection geometry of the network. Signals often traverse multiple intermediaries, meaning that the state perceived by an agent is a non-local function of the network topology newman_2010; barabasi2016. Consequently, errors accumulate along geodesic paths, as seen in network-driven propagation phenomena pastor2001; brockmann2013. The efficiency of information transfer therefore depends critically on the distribution of path lengths and on the small-world properties of the network watts_nature_1998. When information is represented as a real-valued field, biased errors introduce additional randomness, implying that a complete description must incorporate both intrinsic fluctuations and topological constraints.
@@
Applying the framework to canonical network ensembles reveals strong and previously unreported structural effects. Most strikingly, Erdős–Rényi networks exhibit a non-monotonic misinformation profile with two well-separated local maxima, arising from the interplay between the emergence of the giant component, the distribution of geodesic distances, and the formation of short cycles. To our knowledge, this double-peaked behavior has not been documented in prior studies of information distortion on networks. In contrast, scale-free networks suppress misinformation through hub-mediated global coordination, whereas optimally rewired small-world networks suppress it by balancing strong clustering with short characteristic path lengths, consistent with rewiring effects reported in klemn2002. These distinct but complementary mechanisms may offer insights into the reason behind the widespread existence of hub-dominated and small-world architectures in natural and engineered information-processing systems.
@@ II Noisy Information Flow in Biological Networks
Biological networks provide natural examples in which information flow is inherently noisy. This noise might often have important functional consequences. Chemotactic signaling in E. coli fluctuates due to stochastic ligand–receptor binding and downstream processing berg2004coli; tu2008modeling; synaptic transmission is probabilistic and exhibits variability in spike timing faisal2008noise; dayan2001theoretical; gene regulatory networks show intrinsic noise that influences cell-fate decisions elowitz_nature_2000; elowitz2002stochastic; raj2008nature; and immune and protein–interaction networks display variability and crosstalk that reshape collective responses paul2012fundamental; barabasi2004network. These systems highlight that misinformation generated by noise is unavoidable, and amplification, buffering, or redirection of such errors is governed by the structure of the network.
@@
Neuronal networks are another example of path-induced errors. Neurons communicate via action potentials that propagate across synapses, with spike timing and firing rates encoding sensory and cognitive information. Synaptic transmission, however, is inherently noisy as spikes may fail, arrive with temporal jitter, or be distorted by background activities. Such variability might prohibit coding fidelity and lead to pathological synchrony such as epileptic seizures faisal2008noise; dayan2001theoretical; destexhe2003neuronal.
@@
Gene regulatory networks also illustrate how misinformation can arise from intrinsic stochasticity. Transcription factors regulate gene expression through activation and repression, but low copy numbers of mRNA and proteins introduce significant fluctuations. This noise can misguide cell-fate decisions or perturb developmental pathways elowitz2002stochastic; raj2008nature; alon2007introduction. Network motifs such as feed-forward loops and negative-feedback circuits are thought to buffer against such stochasticity.
@@
The immune system offers a further example. Receptor signaling, cytokine communication, and transcriptional responses form a distributed network that are used to discriminate self from non-self. Noisy or ambiguous signals may be amplified inappropriately, generating autoimmunity, while legitimate pathogen cues may fail to elicit a sufficient response paul2012fundamental; perelson1997immune; medzhitov1997; hoffmann2006immune.
@@
Finally, protein–protein interaction (PPI) networks transmit signals via cascades of binding interactions. Non-specific binding, competitive interactions, and pathway crosstalk can misroute signals. The scale-free organization of many PPI networks provides resilience against random failures but also creates vulnerabilities due to reliance on highly connected hubs barabasi2004network; han2004evidence; jeong_nature_2001.
@@
To sum up, the above examples demonstrate that misinformation due to noise is not merely a theoretical construct but a challenge that pervades across biology. They highlight how network topology shapes the reliability of distributed information processing. This reinforce the necessity for a general framework that captures how structures amplify or suppress errors across complex systems, especially across complex networks CLAUSSEN2007; KIM2008.
@@ III.1 Quantifying Misinformation
In reality, the global state of the network is encoded in the random variable $\mathbf{X}$. However, because errors are introduced during propagation, this global information is perceived at the $j$th node as a different random variable, denoted $\mathbf{X}_{j}$. Consequently, the original PDF $P(x)$ is viewed at node $j$ as a modified distribution $Q^{(j)}(x)$. The misinformation encountered at node $j$ can therefore be quantified by measuring the relative distance between the true distribution $P(x)$ and the perceived distribution $Q^{(j)}(x)$. A natural and widely used metric for this purpose is the relative entropy, or Kullback–Leibler (KL) divergence desurvire_2009; cover_2006. The KL divergence between $P(x)$ (true) and $Q^{(j)}(x)$ (perceived) is defined as
@@
where $H(P)$ is the Shannon entropy of $P(x)$ and $H(P,Q^{(j)})$ is the cross-entropy that measures the coding cost of $P(x)$ under $Q^{(j)}(x)$ cover_2006; mackay_2003. (please refer to the right panel of Fig. 1).
@@
where the mean-field coefficients AA and $B$ depend on the distance statistics of the network and satisfy $0<A,B<1$. One can notice that the misinformation measure vanishes in the large-variance limit as the extremal distortion is determined solely by the variance of the Gaussian distribution. This means that when the intrinsic uncertainty of the global state dominates over network-induced biases, the topology is irrelevant. When the true distribution is broad and unstructured, finite perception errors cannot be distinguished, and the network ceases to act as an “information-distorting medium”. As $\sigma\to 0$ in Eq. 10, this bound diverges, implying that the Dirac-delta limit is ill-defined in this coarse-grained formalism and therefore requires a separate treatment.
@@ IV Complete Graph Topology
The true distribution $P(x)$ can be expressed in discrete form as oppenheim_book
@@
Because the KL divergence between identical distributions is zero cover_2006; mackay_2003, we obtain
@@ V Analytical case study of a Dirac-Delta true state
For distance-regular graphs—whose geodesic-shell structure is fully characterized by an intersection array brouwer1989distance—the shell occupations $b^{(j)}_{m}$ are independent of $j$, reducing Eq. 21 to
@@
which depends solely on the coefficients of the intersection array. For asymptotically tree-like networks (e.g., Erdős–Rényi graphs bollobas1998random or configuration models newman2001random), $b_{m}$ admits the closed-form approximation $b_{m}\approx c^{m}$ for mean branching factor $c$, leading to the analytic extremality condition
@@
More generally, since $b^{(j)}_{m}=(A^{m}\mathbf{1})_{j}$, the spectral decomposition of the adjacency operator chung1997spectral yields the closed-form spectral condition
@@
where $\lambda_{k}$ and $\mathbf{v}_{k}$ are the eigenvalues and eigenvectors of AA. Thus, fully explicit optimality criteria arise whenever the geodesic structure of the network can be written in closed form—either through distance-regularity, asymptotic tree-likeness, or spectral symmetry.
@@ VI Numerical Results
These three topologies were chosen because they span distinct geometric and dynamical regimes relevant to stochastic error propagation. Erdős–Rényi graphs represent maximally random connectivity, with approximately Poissonian degree fluctuations erdos_1959; haenggi_ieee_2009. Watts–Strogatz networks interpolate between regular lattices and random graphs, capturing the coexistence of short paths and high clustering that shapes diffusion and communication processes watts_nature_1998; newman_pnas_2001. Barabási–Albert networks produce heterogeneous, hub-dominated structures through preferential attachment, resulting in heavy-tailed degree distributions and nontrivial transport pathways barabasi_science_1999; albert_nature_1999. Because these ensembles differ sharply in geodesic organization, clustering, and degree heterogeneity, they induce markedly different patterns of error accumulation and dissipation. Together, they provide a representative cross-section of the topological regimes in which misinformation dynamics typically exhibit nonlinear amplification, suppression, or saturation behaviors.
@@ VI.1 Erdős–Rényi Networks
For Erdős–Rényi (ER) graphs $\mathcal{G}(n,p)$, we considered networks with $n=5000$ nodes and varied the link probability $p$, giving an average degree $\langle k\rangle=np$ erdos_1959; bollobas_2001. Although few real-world systems follow a pure ER structure, the model is widely used as a null ensemble when links can be approximated as forming independently with equal probability pperdos_1959; bollobas_2001. Such approximations arise in relatively homogeneous settings, including classroom social networks wasserman_1994, local peer-to-peer wireless setups li_wcmc_2003, and laboratory collaboration groups newman_pnas_2001. Similarly, locally dense neuronal microcircuits exhibit partially random synaptic connectivity with experimentally observed connection probabilities of $5\%$$30\%$song_plos_2005; lefort_neuron_2009; perin_pnas_2011. For networks of size $n\sim 10^{3}$, these values correspond to link probabilities $p\approx 0.05$$0.3$, yielding mean degrees in the range $\langle k\rangle\sim 50$300300, consistent with empirical densities. Although real systems also display clustering and spatial constraints, they are typically sparse with roughly probabilistic links, making the ER model a natural and widely used baseline for comparison.
Figure 2 reveals a striking feature of the Erdős–Rényi network—the normalized misinformation exhibits a strongly non-monotonic dependence on the link probability, with two distinct local maxima appearing in the sparse regime. For the sparsest connected graphs, the misinformation begins at a high value and then drops rapidly to a shallow minimum just above the connectivity threshold, where the network remains almost tree-like and supports only a single long path between most node pairs, limiting the accumulation of errors. As the link probability increases further, additional edges shorten some geodesics while still preserving many long routes; this coexistence leads to a second, smaller maximum at slightly higher connectivity, reflecting amplification of path-accumulated binomial errors. Beyond this regime, the Erdős–Rényi network becomes progressively more connected, and the average shortest path collapses toward its logarithmic limit chung_lu2002. This causes the misinformation to decrease monotonically and eventually approach zero when the network is fully connected. Thus, the largest distortions arise not at high connectivity but within a specific sparse range where information spreads widely yet still traverses error-prone long paths. This result is a structural feature emerging naturally from our model and, to our knowledge, not previously reported.
@@ VI.2 Watts–Strogatz Networks
While Erdős–Rényi graphs provide a useful baseline, their lack of clustering and structural heterogeneity limits their relevance for most real-world systems. Empirical studies across social, biological, and technological domains consistently show high clustering, modular organization, heterogeneous degree distributions, and short characteristic path lengths watts_nature_1998; barabasi_science_1999; newman_siam_2003; albert_rmp_2002; girvan_pnas_2002. These features fall outside the ER framework and motivate the use of more structured generative models. The Watts–Strogatz (WS) model watts_nature_1998 captures the emergence of small-world structure by randomly rewiring edges in a regular lattice with probability $\beta$. This parameter controls the balance between local clustering and global integration: $\beta=0$ produces a highly clustered ring lattice with long path lengths, increasing $\beta$ introduces shortcuts that rapidly reduce these lengths yet largely preserving local neighborhoods, and $\beta=1$ corresponds to a random graph with very low clustering.
Small-world patterns are widely observed across biological systems, although the underlying networks are not explicitly formed by rewiring. Neural connectomes, protein–interaction and gene-regulatory networks, metabolic pathways, and ecological food webs exhibit high clustering paired with short paths jeong_nature_2001; han2004evidence; albert_jcell_2005; ravasz_science_2002; barabasi2004network; montoya_jtb_2002; amaral_pnas_2000. Comparisons with WS reference networks suggest that many biological systems occupy an intermediate effective rewiring regime, typically $\beta\approx 0.01$$0.2$ latora2001efficient; sporns2004small; humphries2008network; kaiser2004spatial. Table 1 lists effective rewiring values inferred from such empirical comparisons.
@@
Within this transition, an intermediate rewiring level forms an effective critical regime: local neighborhoods remain coherent, while a modest number of shortcuts provide efficient long-range coordination and suppress error accumulation watts_nature_1998; strogatz2001. This crossover from diffusion-dominated to shortcut-mediated transport mirrors phenomena in disordered media and complex networks barrat2008; dorogovtsev2008. Consequently, the minimum misinformation arises from the interplay between local structure (set by nearest neighbors) and the partial randomness introduced by rewiring, a balance that supports robust and efficient information flow in many natural and engineered systems.
@@
Table 1: Approximate effective rewiring probabilities ($\beta$) for representative biological networks, obtained by matching empirical characteristic path lengths ($L$) to Watts–Strogatz predictions newman_pre_1999; newman_prl_2000. The resulting interpretations show how biological systems occupy ordered, intermediate, or random topological regimes—mirroring the misinformation behaviors observed in our simulations.
Figure 4: Normalized misinformation ($I_{M}$) versus the number of edges added per new node ($m$) in Barabási–Albert networks with $n=5000$ nodes and a $\delta$-distributed true state. Misinformation initially decreases monotonically with increasing $m$, reflecting the well-known role of hubs in enhancing robustness and stabilizing information flow in scale-free structures barabasi_science_1999; albert2000error; barrat2008. Beyond an intermediate threshold, however, misinformation increases again, indicating a regime in which hub-driven fluctuations accumulate and propagate through the network. In the extremely sparse limit (small $m$), BA networks outperform random and small-world networks of comparable density, owing to their highly connected hubs, which act as efficient integrators and provide redundant communication pathways. Shaded regions show the regimes occupied by real-world scale-free networks and their corresponding misinformation levels, based on the comparable connectivity estimates reported in Table 2.
@@
Metabolic network (E. coli) ATP, NADH, H2O, CO2 Hub metabolites act as universal biochemical carriers that integrate multiple pathways, stabilize global flux balance, and buffer stochastic fluctuations arising from local reaction noise jeong2000. 2.2–2.4 3–5
@@
Protein–protein interaction (PPI) network Kinases, chaperones Highly connected proteins integrate signaling modules, buffer proteomic perturbations, stabilize folding environments, and prevent miscommunication within interaction pathways jeong_nature_2001. 2.4–2.7 3–6
Neuronal connectome Thalamus, precuneus, rich-club hubs High-degree cortical hubs support long-range integration, enable rapid routing between distributed brain modules, and mitigate noise accumulation in multi-step communication processes eguiluz2005scale. 2.0–2.3 20–40
Table 2: Representative biological scale-free networks with example hubs, functional roles relevant to misinformation suppression, and their characteristic degree-distribution exponents $\gamma$ and approximate Barabási–Albert parameter $m$ ($\langle k\rangle\approx 2m$). Neuronal ranges vary due to differences in connectome resolution, parcellation, and thresholding across datasets bullmore2009; hagmann2008mapping; modha2010network; van2010exploring.

### v3

# arXiv:2512.19465v3[physics.soc-ph] 02 Mar 2026

@@ I Introduction
Quantification and dynamics of misinformation have been studied across journalism, sociology, neuroscience, collective biological systems, distributed computing, and networked cyber-physical systems, but a unified mathematical understanding remains limited. Prior studies have offered domain-specific or simulation-driven insights, such as qualitative analyses of error propagation based on connectivity measures Krol and Kukla (2012), error accumulation in deep neural architectures Li et al. (2017), and preferred interaction topologies in sociotechnical and biological settings Klemm and Bornholdt (2005). Efforts to restrict misinformation on online platforms Ghoshal et al. (2012) and to characterize the spread of true versus false information in complex networks Nian et al. (2021) provide valuable perspectives, but do not yield a general analytical framework for quantifying misinformation in distributed systems. Complex networks display emergent macroscopic behavior and are vulnerable to localized failures and perturbations that can propagate through the interconnected structure, leading to dysfunction. Recent theoretical and computational frameworks characterize these effects in terms of robustness and resilience, thus providing tools to analyze, predict, and mitigate network breakdown Ji et al. (2023); Artime et al. (2024).
Information flow in any multi-agent system is fundamentally shaped by connection geometry of the network. Signals often traverse multiple intermediaries, meaning that the state perceived by an agent is a non-local function of the network topology Newman (2010); Barabási and Pósfai (2016). Consequently, errors accumulate along geodesic paths, as seen in network-driven propagation phenomena Pastor-Satorras and Vespignani (2001); Brockmann and Helbing (2013). The efficiency of information transfer therefore depends critically on the distribution of path lengths and on the small-world properties of the network Watts and Strogatz (1998). When information is represented as a real-valued field, biased errors introduce additional randomness, implying that a complete description must incorporate both intrinsic fluctuations and topological constraints.
@@
Applying the framework to canonical network ensembles reveals strong and previously unreported structural effects. Most strikingly, Erdős–Rényi networks exhibit a non-monotonic misinformation profile with two well-separated local maxima, arising from the interplay between the emergence of the giant component, the distribution of geodesic distances, and the formation of short cycles. To our knowledge, this double-peaked behavior has not been documented in prior studies of information distortion on networks. In contrast, scale-free networks suppress misinformation through hub-mediated global coordination, whereas optimally rewired small-world networks suppress it by balancing strong clustering with short characteristic path lengths, consistent with rewiring effects reported in Klemm and Eguíluz (2002). These distinct but complementary mechanisms may offer insights into the reason behind the widespread existence of hub-dominated and small-world architectures in natural and engineered information-processing systems.
@@ II Noisy Information Flow in Biological Networks
Biological networks provide natural examples in which information flow is inherently noisy. This noise might often have important functional consequences. Chemotactic signaling in E. coli fluctuates due to stochastic ligand–receptor binding and downstream processing Berg (2004); Tu (2008); synaptic transmission is probabilistic and exhibits variability in spike timing Faisal et al. (2008); Dayan and Abbott (2001); gene regulatory networks show intrinsic noise that influences cell-fate decisions Elowitz and Leibler (2000); Elowitz et al. (2002); Raj and van Oudenaarden (2008); and immune and protein–interaction networks display variability and crosstalk that reshape collective responses Paul (2012); Barabási and Oltvai (2004). These systems highlight that misinformation generated by noise is unavoidable, and amplification, buffering, or redirection of such errors is governed by the structure of the network.
@@
Neuronal networks are another example of path-induced errors. Neurons communicate via action potentials that propagate across synapses, with spike timing and firing rates encoding sensory and cognitive information. Synaptic transmission, however, is inherently noisy as spikes may fail, arrive with temporal jitter, or be distorted by background activities. Such variability might prohibit coding fidelity and lead to pathological synchrony such as epileptic seizures Faisal et al. (2008); Dayan and Abbott (2001); Destexhe and Paré (1999).
@@
Gene regulatory networks also illustrate how misinformation can arise from intrinsic stochasticity. Transcription factors regulate gene expression through activation and repression, but low copy numbers of mRNA and proteins introduce significant fluctuations. This noise can misguide cell-fate decisions or perturb developmental pathways Elowitz et al. (2002); Raj and van Oudenaarden (2008); Alon (2007). Network motifs such as feed-forward loops and negative-feedback circuits are thought to buffer against such stochasticity.
@@
The immune system offers a further example. Receptor signaling, cytokine communication, and transcriptional responses form a distributed network that are used to discriminate self from non-self. Noisy or ambiguous signals may be amplified inappropriately, generating autoimmunity, while legitimate pathogen cues may fail to elicit a sufficient response Paul (2012); Perelson and Weisbuch (1997); Medzhitov and Janeway (1997); Hoffmann et al. (2002).
@@
Finally, protein–protein interaction (PPI) networks transmit signals via cascades of binding interactions. Non-specific binding, competitive interactions, and pathway crosstalk can misroute signals. The scale-free organization of many PPI networks provides resilience against random failures but also creates vulnerabilities due to reliance on highly connected hubs Barabási and Oltvai (2004); Han et al. (2004); Jeong et al. (2001).
@@
To sum up, the above examples demonstrate that misinformation due to noise is not merely a theoretical construct but a challenge that pervades across biology. They highlight how network topology shapes the reliability of distributed information processing. This reinforce the necessity for a general framework that captures how structures amplify or suppress errors across complex systems, especially across complex networks Claussen (2007); Kim and Wilhelm (2008).
@@ III.1 Quantifying Misinformation
In reality, the global state of the network is encoded in the random variable $\mathbf{X}$. However, because errors are introduced during propagation, this global information is perceived at the $j$th node as a different random variable, denoted $\mathbf{X}_{j}$. Consequently, the original PDF $P(x)$ is viewed at node $j$ as a modified distribution $Q^{(j)}(x)$. The misinformation encountered at node $j$ can therefore be quantified by measuring the relative distance between the true distribution $P(x)$ and the perceived distribution $Q^{(j)}(x)$. A natural and widely used metric for this purpose is the relative entropy, or Kullback–Leibler (KL) divergence Desurvire (2009); Cover and Thomas (2006). The KL divergence between $P(x)$ (true) and $Q^{(j)}(x)$ (perceived) is defined as
@@
where $H(P)$ is the Shannon entropy of $P(x)$ and $H(P,Q^{(j)})$ is the cross-entropy that measures the coding cost of $P(x)$ under $Q^{(j)}(x)$ Cover and Thomas (2006); MacKay (2002). (please refer to the right panel of Fig. 1).
@@
where the mean-field coefficients $A$ and $B$ depend on the distance statistics of the network and satisfy $0<A,B<1$. One can notice that the misinformation measure vanishes in the large-variance limit as the extremal distortion is determined solely by the variance of the Gaussian distribution. This means that when the intrinsic uncertainty of the global state dominates over network-induced biases, the topology is irrelevant. When the true distribution is broad and unstructured, finite perception errors cannot be distinguished, and the network ceases to act as an “information-distorting medium”. As $\sigma\to 0$ in Eq. 10, this bound diverges, implying that the Dirac-delta limit is ill-defined in this coarse-grained formalism and therefore requires a separate treatment.
@@ IV Complete Graph Topology
The true distribution $P(x)$ can be expressed in discrete form as Oppenheim et al. (1999)
@@
Because the KL divergence between identical distributions is zero Cover and Thomas (2006); MacKay (2002), we obtain
@@ V Analytical case study of a Dirac-Delta true state
For distance-regular graphs—whose geodesic-shell structure is fully characterized by an intersection array Brouwer et al. (1989)—the shell occupations $b^{(j)}_{m}$ are independent of $j$, reducing Eq. 21 to
@@
which depends solely on the coefficients of the intersection array. For asymptotically tree-like networks (e.g., Erdős–Rényi graphs Bollobás (1998) or configuration models Newman et al. (2001)), $b_{m}$ admits the closed-form approximation $b_{m}\approx c^{m}$ for mean branching factor $c$, leading to the analytic extremality condition
@@
More generally, since $b^{(j)}_{m}=(A^{m}\mathbf{1})_{j}$, the spectral decomposition of the adjacency operator Chung (1997) yields the closed-form spectral condition
@@
where $\lambda_{k}$ and $\mathbf{v}_{k}$ are the eigenvalues and eigenvectors of $A$. Thus, fully explicit optimality criteria arise whenever the geodesic structure of the network can be written in closed form—either through distance-regularity, asymptotic tree-likeness, or spectral symmetry.
@@ VI Numerical Results
These three topologies were chosen because they span distinct geometric and dynamical regimes relevant to stochastic error propagation. Erdős–Rényi graphs represent maximally random connectivity, with approximately Poissonian degree fluctuations Erdős and Rényi (1959); Haenggi et al. (2009). Watts–Strogatz networks interpolate between regular lattices and random graphs, capturing the coexistence of short paths and high clustering that shapes diffusion and communication processes Watts and Strogatz (1998); Newman (2001). Barabási–Albert networks produce heterogeneous, hub-dominated structures through preferential attachment, resulting in heavy-tailed degree distributions and nontrivial transport pathways Barabási and Albert (1999); Albert et al. (1999). Because these ensembles differ sharply in geodesic organization, clustering, and degree heterogeneity, they induce markedly different patterns of error accumulation and dissipation. Together, they provide a representative cross-section of the topological regimes in which misinformation dynamics typically exhibit nonlinear amplification, suppression, or saturation behaviors.
@@ VI.1 Erdős–Rényi Networks
For Erdős–Rényi (ER) graphs $\mathcal{G}(n,p)$, we considered networks with $n=5000$ nodes and varied the link probability $p$, giving an average degree $\langle k\rangle=np$ Erdős and Rényi (1959); Bollobás (2001). Although few real-world systems follow a pure ER structure, the model is widely used as a null ensemble when links can be approximated as forming independently with equal probability ppErdős and Rényi (1959); Bollobás (2001). Such approximations arise in relatively homogeneous settings, including classroom social networks Wasserman and Faust (1994), local peer-to-peer wireless setups Li (2003), and laboratory collaboration groups Newman (2001). Similarly, locally dense neuronal microcircuits exhibit partially random synaptic connectivity with experimentally observed connection probabilities of $5\%$$30\%$Song et al. (2005); Lefort et al. (2009); Perin et al. (2011). For networks of size $n\sim 10^{3}$, these values correspond to link probabilities $p\approx 0.05$$0.3$, yielding mean degrees in the range $\langle k\rangle\sim 50$300300, consistent with empirical densities. Although real systems also display clustering and spatial constraints, they are typically sparse with roughly probabilistic links, making the ER model a natural and widely used baseline for comparison.
Figure 2 reveals a striking feature of the Erdős–Rényi network—the normalized misinformation exhibits a strongly non-monotonic dependence on the link probability, with two distinct local maxima appearing in the sparse regime. For the sparsest connected graphs, the misinformation begins at a high value and then drops rapidly to a shallow minimum just above the connectivity threshold, where the network remains almost tree-like and supports only a single long path between most node pairs, limiting the accumulation of errors. As the link probability increases further, additional edges shorten some geodesics while still preserving many long routes; this coexistence leads to a second, smaller maximum at slightly higher connectivity, reflecting amplification of path-accumulated binomial errors. Beyond this regime, the Erdős–Rényi network becomes progressively more connected, and the average shortest path collapses toward its logarithmic limit Chung and Lu (2002). This causes the misinformation to decrease monotonically and eventually approach zero when the network is fully connected. Thus, the largest distortions arise not at high connectivity but within a specific sparse range where information spreads widely yet still traverses error-prone long paths. This result is a structural feature emerging naturally from our model and, to our knowledge, not previously reported.
@@ VI.2 Watts–Strogatz Networks
While Erdős–Rényi graphs provide a useful baseline, their lack of clustering and structural heterogeneity limits their relevance for most real-world systems. Empirical studies across social, biological, and technological domains consistently show high clustering, modular organization, heterogeneous degree distributions, and short characteristic path lengths Watts and Strogatz (1998); Barabási and Albert (1999); Newman (2003); Albert and Barabási (2002); Girvan and Newman (2002). These features fall outside the ER framework and motivate the use of more structured generative models. The Watts–Strogatz (WS) model Watts and Strogatz (1998) captures the emergence of small-world structure by randomly rewiring edges in a regular lattice with probability $\beta$. This parameter controls the balance between local clustering and global integration: $\beta=0$ produces a highly clustered ring lattice with long path lengths, increasing $\beta$ introduces shortcuts that rapidly reduce these lengths yet largely preserving local neighborhoods, and $\beta=1$ corresponds to a random graph with very low clustering.
Small-world patterns are widely observed across biological systems, although the underlying networks are not explicitly formed by rewiring. Neural connectomes, protein–interaction and gene-regulatory networks, metabolic pathways, and ecological food webs exhibit high clustering paired with short paths Jeong et al. (2001); Han et al. (2004); Albert (2005); Ravasz et al. (2002); Barabási and Oltvai (2004); Montoya and Solé (2002); Amaral et al. (2000). Comparisons with WS reference networks suggest that many biological systems occupy an intermediate effective rewiring regime, typically $\beta\approx 0.01$$0.2$ Latora and Marchiori (2001); Sporns and Zwi (2004); Humphries and Gurney (2008); Kaiser and Hilgetag (2004). Table 1 lists effective rewiring values inferred from such empirical comparisons.
@@
Within this transition, an intermediate rewiring level forms an effective critical regime: local neighborhoods remain coherent, while a modest number of shortcuts provide efficient long-range coordination and suppress error accumulation Watts and Strogatz (1998); Strogatz (2001). This crossover from diffusion-dominated to shortcut-mediated transport mirrors phenomena in disordered media and complex networks Barrat et al. (2008); Dorogovtsev et al. (2008). Consequently, the minimum misinformation arises from the interplay between local structure (set by nearest neighbors) and the partial randomness introduced by rewiring, a balance that supports robust and efficient information flow in many natural and engineered systems.
@@
Table 1: Approximate effective rewiring probabilities ($\beta$) for representative biological networks, obtained by matching empirical characteristic path lengths ($L$) to Watts–Strogatz predictions Newman and Watts (1999); Newman et al. (2000). The resulting interpretations show how biological systems occupy ordered, intermediate, or random topological regimes—mirroring the misinformation behaviors observed in our simulations.
Figure 4: Normalized misinformation ($I_{M}$) versus the number of edges added per new node ($m$) in Barabási–Albert networks with $n=5000$ nodes and a $\delta$-distributed true state. Misinformation initially decreases monotonically with increasing $m$, reflecting the well-known role of hubs in enhancing robustness and stabilizing information flow in scale-free structures Barabási and Albert (1999); Albert et al. (2000); Barrat et al. (2008). Beyond an intermediate threshold, however, misinformation increases again, indicating a regime in which hub-driven fluctuations accumulate and propagate through the network. In the extremely sparse limit (small $m$), BA networks outperform random and small-world networks of comparable density, owing to their highly connected hubs, which act as efficient integrators and provide redundant communication pathways. Shaded regions show the regimes occupied by real-world scale-free networks and their corresponding misinformation levels, based on the comparable connectivity estimates reported in Table 2.
@@
Metabolic network (E. coli) ATP, NADH, H2O, CO2 Hub metabolites act as universal biochemical carriers that integrate multiple pathways, stabilize global flux balance, and buffer stochastic fluctuations arising from local reaction noise Jeong et al. (2000). 2.2–2.4 3–5
@@
Protein–protein interaction (PPI) network Kinases, chaperones Highly connected proteins integrate signaling modules, buffer proteomic perturbations, stabilize folding environments, and prevent miscommunication within interaction pathways Jeong et al. (2001). 2.4–2.7 3–6
Neuronal connectome Thalamus, precuneus, rich-club hubs High-degree cortical hubs support long-range integration, enable rapid routing between distributed brain modules, and mitigate noise accumulation in multi-step communication processes Eguíluz et al. (2005). 2.0–2.3 20–40
Table 2: Representative biological scale-free networks with example hubs, functional roles relevant to misinformation suppression, and their characteristic degree-distribution exponents $\gamma$ and approximate Barabási–Albert parameter $m$ ($\langle k\rangle\approx 2m$). Neuronal ranges vary due to differences in connectome resolution, parcellation, and thresholding across datasets Bullmore and Sporns (2009); Hagmann et al. (2008); Modha and Singh (2010); van den Heuvel and Hulshoff Pol (2010).